Dunder-Mifflin/Saber has announced that it will begin paying dividends with the first dividend of $3.50 per share to be paid in three years and grow at 6% per year afterward. If you require a return of 18%, what is the most you would pay for the stock today?
Answer #1

Value after year 3=(D3*Growth rate)/(Required rate-Growth rate)

=(3.5*1.06)/(0.18-0.06)

=(3.5*1.06)/(0.18-0.06)

=$30.9167

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.5/1.18^3+30.9167/1.18^3

=$20.95(Approx).